OpenStudy (anonymous):

The binomials shown below will be multiplied to produce a quadratic trinomial. (x + p)(x + q) Which part of the trinomial will equal the sum of p and q? a.the coefficient of the x term b.the constant term c.the degree d.the coefficient of the x^2 term

OpenStudy (anonymous):

Multiply the two binomials using FOIL. \((x + p)(x + q) \) \(= x^2 + qx + px + pq\) \(=x^2 + (p + q)x + pq\)

OpenStudy (anonymous):

So what would the answer be? I don't understand @MaxwellSmart

OpenStudy (anonymous):

Do you see anywhere the sum of p and q, such as p + q?

OpenStudy (anonymous):

What does (p + q) multiply?

OpenStudy (anonymous):

I don't know I'm really confused....

OpenStudy (anonymous):

I break down the question in your problem into small questions that are easy to answer. You don't need to know much to answer my questions. I asked if you see (p + q) in what I did. Do you see it or not?

OpenStudy (anonymous):

you multiplied it by x?

OpenStudy (anonymous):

Exactly. You do see (p + q) in the last equation I wrote. It is multiplying the x in the middle term. The number that multiplies a variable is called coefficient. That means (p + q) is the coefficient of the x-term.

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

You're welcome.